### v1

# Constacyclic codes with best-known parameters

Abstract
In this paper, we construct several infinite families of $q$-ary constacyclic codes over a finite field $\mathbb{F}_{q}$ with length $n$, dimension around $n/2$, and minimum distance at least $cn/\log_{q}n$ for some positive constant $c$. They contain many constacyclic codes with optimal, or almost-optimal, or best-known parameters. We also consider various forms of the length $n$.
Z. Chen’s research was supported by the Scientific Research Innovation Project of Graduate School of South China Normal University. M. Sha’s research was supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515010635). Zekai Chen and Min Sha are with School of Mathematical Sciences, South China Normal University, Guangzhou, Guangdong, 510631, China. (email: chenzk@m.scnu.edu.cn, min.sha@m.scnu.edu.cn)
1 Introduction
1.1 Background
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$. A $q$-ary $[n,k,d]$ linear code $\mathcal{C}$ is a $k$-dimensional linear subspace of $\mathbb{F}_{q}^{n}$ with minimum distance $d$. Here, $n$ is called the length of $\mathcal{C}$. The linear code $\mathcal{C}$ is said to be distance optimal if there is no $q$-ary $n,k,\geq d+1]$ linear code. In addition, $\mathcal{C}$ is said to be distance almost-optimal if there is a $q$-ary $[n,k,d+1]$ linear code but there is no $q$-ary $n,k,\geq d+2]$ linear code. Moreover, $\mathcal{C}$ is said to be distance best-known if $d$ is the maximal minimum distance of all known $q$-ary linear codes with length $n$ and dimension $k$.
Let $\lambda\in\mathbb{F}_{q}^{*}$. The linear code $\mathcal{C}$ is called a $\lambda$-constacyclic code if $(\lambda c_{n-1},c_{0},\ldots,c_{n-2})\in\mathcal{C}$ for any $(c_{0},c_{1},\ldots,c_{n-1})\in\mathcal{C}$. Define the map
$\phi:\,\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$
$(c_{0},c_{1},\ldots,c_{n-1})\mapsto c_{0}+c_{1}x+\cdots+c_{n-1}x^{n-1}.$
It is well-known that every ideal of the residue class ring $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$ is principal. Then, the linear code $\mathcal{C}$ is called a $\lambda$-constacyclic code if and only if $\phi(\mathcal{C})$ is an ideal of the ring $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$. So, we identify $\mathcal{C}$ with $\phi(\mathcal{C})$ for any $\lambda$-constacyclic code $\mathcal{C}$. Let $\mathcal{C}=\langle g(x)\rangle$ be a $q$-ary $\lambda$-constacyclic code of length $n$, where $g(x)$ is monic and has the smallest degree among all the non-zero codewords of $\mathcal{C}$. Then, $g(x)$ is a factor of $x^{n}-\lambda$ and is called the generator polynomial of $\mathcal{C}$.
In particular, when $\lambda=1$, $\lambda$-constacyclic codes are classical cyclic codes; and when $\lambda=-1$ and $q$ is odd, $\lambda$-constacyclic codes are negacyclic codes. We refer to [16, 24, 27] for more details about linear codes and cyclic codes.
Constacyclic codes have attracted extensive attentions due to their performance and applications. For example, several new constacyclic codes that improve the minimum distance of currently best known linear codes were found in [1, 2], several infinite classes of MDS constacyclic codes were constructed in [14, 18, 36], and several infinite families of distance-optimal constacyclic codes were constructed in [12, 42]. In addition, constacyclic codes have important applications in constructing other kinds of codes, such as locally repairable codes [3], symbol-pair codes [4], and quantum codes [5, 17]. We refer to [30] for more details and references about constacyclic codes.
1.2 Related works
It is a long-standing open problem in coding theory that whether there exists an asymptotically good sequence of cyclic codes (see [25]). Therefore, it is interesting to construct infinite families of cyclic, negacyclic or constacyclic codes whose dimensions and minimum distances are both large.
In recent years, researchers pay much attention in constructing infinite families of such codes with parameters $[n,k,d]$ such that $k$ is around $n/2$ and $d$ is at least $c\sqrt{n}$ for some positive constant $c$ depending only on $q$ (this lower bound is called the square-root-like lower bound).
Some binary quadratic residue codes with parameters $[n,(n+1)/2,d]$ and their even-weight subcodes with parameters [n,(n1)/2,d+1][n,(n-1)/2,d+1], where $d\geq\sqrt{n}$ and $n$ is a prime with $n\equiv\pm 1\pmod{8}$, were constructed in [16, Section 6.6]. Two infinite families of binary cyclic codes with length $2^{m}-1$ and dimension near $2^{m-1}$ and with a square-root-like lower bound on minimum distance were constructed in [34]. Recently, the work in [34] was extended to the 4-ary case in [28] and the $2^{s}$-ary case in [19]. We refer to [20, 21, 22, 29, 31] for more infinite families of binary cyclic codes with a square-root-like lower bound, [10, 22] for infinite families of ternary cyclic codes with a square-root-like lower bound, and [6, 7, 33, 40, 41] for infinite families of $q$-ary cyclic codes with a square-root-like lower bound. In addition, several infinite families of $q$-ary negacyclic codes and constacyclic codes were constructed in [11, 30, 38] with a square-root-like lower bound.
In particular, in [32] the first infinite family of binary $[n,(n+1)/2,d]$ cyclic codes was constructed with minimum distance $d$ provably much better than the square-root-like bound, where $n=2^{p}-1$, $p$ is an odd prime and $d\geq cn/\log_{2}n$ for some positive constant $c$. Roughly speaking, the defining set of these codes is the union of the first half of cyclotomic cosets with size $p$.
Motivated by [32], several works have been done recently for constructing more $q$-ary linear codes with a lower bound like $cn/\log_{q}n$ on their minimum distances. We summarize them as follows:
  • In [8], one infinite family of $q$-ary negacyclic $n,(n+1)/2,\geq cn/\log_{q}n]$ codes with $q\equiv 3\pmod{4}$ and $n=\frac{q^{p}-1}{q-1}$ was constructed, and one infinite family of $q$-ary negacyclic $n,n/2,\geq cn/\log_{q}n]$ codes with $q\equiv 1\pmod{4}$ and $n=\frac{q^{p}-1}{2}$ was also constructed.
  • In [9], two infinite families of repeated-root binary cyclic codes with parameters $2n,k,\geq(n-1)/\log_{2}n]$, where $n=2^{p}-1$ and $k\geq n$, were constructed.
  • In [20], two infinite families of binary $n,(n+1)/2,\geq cn/\log_{2}n]$ cyclic codes with $n=2^{p^{2}}-1$ or $2^{p_{1}p_{2}}-1$ were construced in [20], where $p_{1}$ and $p_{2}$ are two distinct primes.
  • In [23], an improvement of the lower bound on minimum distance in [32] was given, and an infinite family of ternary $n,(n+2)/2,\geq cn/\log_{3}n]$ cyclic codes was constructed, where $n=3^{p}-1$.
  • In [31], an infinite family of binary $2^{m}-1,k,\geq\lfloor 2^{m-1}/m\rfloor+2]$ cyclic codes was constructed for any integer $m\geq 3$, where $k\geq 2^{m-1}-1$.
Most recently, an infinite family of binary BCH codes was constructed in [39] with length $n=(2^{p_{1}}-1)\cdots(2^{p_{s}}-1)$, dimension at least $(n+1)/2$, and minimum distance at least $\lceil(n-1)/(p_{1}\cdots p_{s})\rceil$, where $p_{1},\ldots,p_{s}$ are distinct primes. We also refer to [26] for BCH codes with minimum distances propotional to code lengths.
1.3 Our contributions
In this paper, motivated by [32] and generalizing the constructions in [8, 9, 20, 23, 31, 32], we propose the following general construction for $q$-ary constacyclic codes: first we classify the cyclotomic cosets according to their sizes (that is, two cosets are in the same class if and only if they have the same size), and then we take the defining set as the union of the first half of cosets in each class (we will use the ceiling function and the floor function respectively to achieve “the first half”).
In order to obtain good lower bounds for minimum distance, in this paper we focus on the case when there are exactly two kinds of cyclotomic cosets. For the minimum distance, if $N$ is the number of cyclotomic cosets with maximum size, by construction one may guess that the first $\lfloor N/2\rfloor$ consecutive elements are in the defining set. However, this does not always hold.
The main contributions of this paper are as follows:
  • We establish the lower bound $\lfloor\frac{qN}{2(q-1)}\rfloor$ on minimum distances for $q$-ary constacyclic codes constructed by using the ceiling function, and the same lower bound in many situations for those codes constructed by using the floor function (see Theorem 3.1). This improves the lower bounds on minimum distance in [8, Theorems 10, 15 and 16] for some infinite families of $q$-ary negacyclic codes mentioned above. Moreover, for some special cases, we obtain better lower bounds on minimum distances (see Theorem 3.3).
  • We consider various forms of the length $n$: $n=p$, or $\frac{q^{p}-1}{rs}$, or $\Phi_{p^{b}}(q)$, or $\Phi_{p_{1}p_{2}}(q)$, where $p$ is a prime, $r$ is the multiplicative order of $\lambda$, $s$ and $b$ are two positive integers, $p_{1}$ and $p_{2}$ are two distinct primes, and $\Phi_{m}(x)$ is the $m$-th cyclotomic polynomial for any integer $m\geq 1$. We obtain many constacyclic codes with optimal, or almost-optimal, or best-known parameters (according to [15]).
The rest of this paper is organized as follows. In Section 2, we present some preliminary results. In Section 3, we present some general constructions for infinite families of constacyclic codes. In Section 4, we construct two infinite families of constacyclic codes with prime length. In Section 5, we construct two infinite families of constacyclic codes with length $n=\frac{q^{p}-1}{rs}$. In Section 6, we construct several infinite families of constacyclic codes with length $n=\Phi_{p^{b}}(q)$. In Section 7, we construct two infinite families of constacyclic codes with length $n=\Phi_{p_{1}p_{2}}(q)$.
2 Preliminaries
2.1 Cyclotomic cosets
Let $q$ be a power of prime, $n$ be a positive integer with $\gcd(n,q)=1$, and $r\mid q-1$. Let
$\mathbb{Z}_{nr}=\{1,2,\ldots,nr\}$
be the ring of integers modulo $nr$. Here we emphasize that the first element in $\mathbb{Z}_{nr}$ is 1 but not 0, and we need this to unify the statements about cyclic codes and constacyclic codes.
For any $i\in\mathbb{Z}_{nr}$, the $q$-cyclotomic coset of $i$ modulo $nr$ is defined by
$C_{i}^{(q,nr)}=\{iq^{j}\mod nr:\,0\leq j\leq l_{i}-1\}\subseteq\mathbb{Z}_{nr},$
where $l_{i}$ is the least positive integer such that $iq^{l_{i}}\equiv i\mod nr$, that is,
$l_{i}=|C_{i}^{(q,nr)}|.$
The smallest integer in $C_{i}^{(q,nr)}$ is called the coset leader of $C_{i}^{(q,nr)}$. Let $\Gamma_{(q,n,r)}$ be the set of all the coset leaders,
$\Gamma_{(q,n,r)}^{(1)}=\{i\in\Gamma_{(q,n,r)}:\,i\equiv 1\pmod{r}\}$,
and
$Z_{n,r}=\{1+ir:\,0\leq i\leq n-1\}\subseteq\mathbb{Z}_{nr}.$
Clearly, $Z_{n,1}=\mathbb{Z}_{n}$ if $r=1$ (this corresponds to cyclic codes), and
$Z_{n,r}=\bigcup_{i\in\Gamma_{(q,n,r)}^{(1)}}C_{i}^{(q,nr)}.$
We also remark that the case $r=2$ corresponds to negacyclic codes.
Let $\beta$ be a primitive $nr$-th roots of unity and $\lambda=\beta^{n}$. Then, $\lambda$ is an element in $\mathbb{F}_{q}$ with multiplicative order $r$. For any $i\in\mathbb{Z}_{nr}$, let $M_{\beta^{i}}(x)$ denote the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$, then
$M_{\beta^{i}}(x)=\prod_{j\in C_{i}^{(q,nr)}}(x-\beta^{j})\in\mathbb{F}_{q}[x].$
Besides,
$x^{n}-\lambda=x^{n}-\beta^{n}=\prod_{j\in Z_{n,r}}(x-\beta^{j})=\prod_{i\in\Gamma_{(q,n,r)}^{(1)}}M_{\beta^{i}}(x).$
Then, for a $\lambda$-constacyclic code $\mathcal{C}=\langle g(x)\rangle$ with $g(x)\mid x^{n}-\lambda$, the set
$T(\mathcal{C})=\{i\in Z_{n,r}:\,g(\beta^{i})=0\}$
is called the defining set of $\mathcal{C}$ with respect to $\beta$. In addition, the Bose distance of $\mathcal{C}$ is defined to be the largest number of consecutive elements from $Z_{n,r}$ and contained in $T(\mathcal{C})$ (see [13] for some work about Bose distance).
For any positive integer $m$ with $\gcd(q,m)=1$, let $\mathrm{ord}_{m}(q)$ be the the multiplicative order of $q$ modulo $m$. The result in the following lemma is somehow well-known (for instance, see [35, page 3]).
Lemma 2.1.
For any $i\in Z_{n,r}$, we have $l_{i}\mid\mathrm{ord}_{nr}(q)$. In particular, if we write $nr=(q^{m}-1)/s$ for some positive integers $m,s$, then we have $l_{i}\mid m$ for each $i\in Z_{n,r}$.
For any positive integer $l$, let $N_{l}^{(q,nr)}$ be the number of $q$-cyclotomic cosets modulo $nr$ whose coset leaders are in $\Gamma_{q,n,r}^{(1)}$ and sizes are equal to $l$, that is,
$N_{l}^{(q,nr)}=\Big|\big\{C_{i}^{(q,nr)}:\,|C_{i}^{(q,nr)}|=l,i\in Z_{n,r}\big\}\Big|.$
By Lemma 2.1, if $l\nmid\mathrm{ord}_{nr}(q)$, then $N_{l}^{(q,nr)}=0$. In addition, $N_{l}^{(q,nr)}=0$ if $l\neq l_{i}$ for any $i\in Z_{n,r}$.
In the sequel, for simplicity and without confusion we denote
$N_{l}=N_{l}^{(q,nr)},\quad l\geq 1.$
We recall two lemmas in [35, Lemmas 4 and 5], which are useful for determining $N_{l}$.
Lemma 2.2 ([35]).
For any $i\in Z_{n,r}$ and any positive integer $l$, we have
$l_{i}\mid l\iff\frac{nr}{\gcd(nr,q^{l}-1)}\,\Big|\,i.$
Lemma 2.3 ([35]).
For each $l\mid\mathrm{ord}_{nr}(q)$, we have
$N_{l}=\Big(\sum_{j\mid l}\mu(l/j)\sigma(j)\gcd(n,(q^{j}-1)/r)\Big)/l,$
where
$\sigma(j)=\begin{cases}0,&\gcd(r,\frac{nr}{\gcd(nr,q^{j}-1)})>1,\\ 1,&\gcd(r,\frac{nr}{\gcd(nr,q^{j}-1)})=1.\end{cases}$
From Lemma 2.3, we can get a formula about the number of elements $i\in Z_{n,r}$ with $l_{i}$ dividing a given integer.
Lemma 2.4.
For any $l\mid\mathrm{ord}_{nr}(q)$, we have
$\sum_{j\mid l}jN_{j}=\begin{cases}0,&\gcd(r,\frac{nr}{\gcd(nr,q^{l}-1)})>1,\\ \gcd(n,(q^{l}-1)/r),&\gcd(r,\frac{nr}{\gcd(nr,q^{l}-1)})=1.\end{cases}$
Proof.
By Lemma 2.3, we have
$lN_{l}=\sum_{j\mid l}\mu(l/j)\sigma(j)\gcd(n,(q^{j}-1)/r).$
Using the Möbius inversion formula, we directly have
$\sum_{j\mid l}jN_{j}=\sigma(l)\gcd(n,(q^{l}-1)/r),$
which gives the desired result. ∎
2.2 Coset leaders in $Z_{n,r}$
Now, for any positive integer $l$, if $N_{l}^{(q,nr)}>0$, then in $Z_{n,r}$, we list all the coset leaders (in ascending order) corresponding to those $q$-cyclotomic cosets with $l$ elements modulo $nr$, and we denote the $i$-th one by $\delta_{i}^{(q,nr,l)}$ for $1\leq i\leq N_{l}^{(q,nr)}$.
In the sequel, for simplicity and without confusion we denote
$N_{l}=N_{l}^{(q,nr)},\quad l\geq 1,$
$\delta_{i}^{(l)}=\delta_{i}^{(q,nr,l)},\quad 1\leq i\leq N_{l}.$
We want to get some estimates about the coset leaders $\delta_{i}^{(l)}$ for $1\leq i\leq N_{l}$. For this, we need a preparation.
Lemma 2.5.
For any two positive integers $a$ and $i$, the $i$-th element not disivible by $q$ in the set $\{a,a+r,a+2r,\ldots\}$ ((listed in ascending order)) is at least
$a+\left(\lceil qi/(q-1)\rceil-2\right)r.$
Proof.
Denote the set $S_{a,r}=\{a,a+r,a+2r,\ldots\}$, and we list its elements in ascending order. Since $\gcd(q,r)=1$ (due to $r\mid q-1$), we have that every $q$ consecutive elements in $S_{a,r}$ form a complete residue system modulo $q$.
Assume that the $i$-th element not disivible by $q$ in the set $S_{a,r}$ is a+(j1)ra+(j-1)r for some integer $j$. Then, we have either $i=j-\lfloor j/q\rfloor-1$, or $i=j-\lfloor j/q\rfloor$. If $i=j-\lfloor j/q\rfloor-1$, then we have
$i\leq j-(j/q-(q-1)/q)-1,$
which gives
$j\geq\frac{qi}{q-1}+\frac{1}{q-1}.$
If $i=j-\lfloor j/q\rfloor$, then we have
$i\leq j-(j/q-(q-1)/q),$
which implies
$j\geq\frac{qi}{q-1}-1.$
Hence, noticing that $j$ is an integer, we always have
$j\geq\left\lceil\frac{qi}{q-1}-1\right\rceil=\left\lceil\frac{qi}{q-1}\right\rceil-1.$
This yields the desired result. ∎
Now, we are ready to give a lower bound for $\delta_{i}^{(l)}$ with $1\leq i\leq N_{l}$.
Lemma 2.6.
For any positive integer $l$, if $N_{l}>0$, then for any integer $i$ with $1\leq i\leq N_{l}$, we have
$\delta_{i}^{(l)}$ $\geq\frac{nr(1+(\lceil qi/(q-1)\rceil-2)r)}{\gcd(nr,q^{l}-1)}$
$\geq\frac{nr^{2}(i-1)}{\gcd(nr,q^{l}-1)}+\frac{nr}{\gcd(nr,q^{l}-1)}.$
Proof.
First,by Lemma 2.2 we have
$\frac{nr}{\gcd(nr,q^{l}-1)}\,\Big|\,\delta_{i}^{(l)}.$
For simplicity, we denote $t=nr/\gcd(nr,q^{l}-1)$, then $\delta_{i}^{(l)}$ is a multiple of $t$. Since $r\mid q-1$, we have $t\mid n$.
Besides, since $N_{l}>0$, by Lemma 2.4 we have $\gcd(r,t)=1$. So, every $t$ consecutive elements in $Z_{n,r}$ form a complete residue system modulo $t$. If $at$ is the smallest multiple of $t$ contained in $Z_{n,r}$ for some positive integer $a\leq r$, then the set of all the multiples of $t$ in $Z_{n,r}$ is (noticing $\gcd(r,t)=1$):
$S(t)=\{at,at+rt,at+2rt,\ldots,at+jrt\},$
where $j$ is the largest integer such that $at+jrt\leq 1+(n-1)r$. Clearly, we have
$S(t)=\{at,(a+r)t,(a+2r)t,\ldots,(a+jr)t\}.$
Hence, for our purpose it suffices to estimate the $i$-th element in $S(t)$ (listed in ascending order) which is not disivible by $q$.
Since $\gcd(n,q)=1$ and $t\mid n$, we have $\gcd(t,q)=1$. So, it suffices to estimate the $i$-th element in the set $\{a,a+r,\ldots,a+jr\}$ (listed in ascending order) which is not disivible by $q$. Hence, combining this with Lemma 2.5, we obtain
$\begin{split}\delta_{i}^{(l)}&\geq\left(a+(\lceil qi/(q-1)\rceil-2)r\right)t\\ &\geq\frac{nr(1+(\lceil qi/(q-1)\rceil-2)r)}{\gcd(nr,q^{l}-1)}.\end{split}$
This completes the proof of the first lower bound in the lemma.
The second lower bound in the lemma follows directly from the first one, because
$\lceil qi/(q-1)\rceil-2=i+\lceil i/(q-1)\rceil-2\geq i+1-2=i-1.$
We remark that the first lower bound in Lemma 2.6 is somehow optimal for the general case. For example, choosing $q=3,n=80,r=1$, we have $\delta_{2}^{(2)}=20$, which coincides with the lower bounds in Lemma 2.6 (with $i=2,l=2$).
We also need to estimate $\delta_{i}^{(l)}$ further for some special $i$’s.
Lemma 2.7.
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Then, if $N_{m}\geq 1$, we have
$\delta^{(m)}_{\lceil N_{m}/2\rceil+1}\geq\delta^{(m)}_{\lfloor N_{m}/2\rfloor+1}>1+\left(\frac{qN_{m}}{2(q-1)}-2\right)r.$
Proof.
If $N_{m}\geq 1$, by Lemma 2.6 we get
$\delta^{(m)}_{\lceil N_{m}/2\rceil+1}\geq\delta^{(m)}_{\lfloor N_{m}/2\rfloor+1}$ $\geq\frac{nr(1+(\lceil q(\lfloor N_{m}/2\rfloor+1)/(q-1)\rceil-2)r)}{\gcd(nr,q^{m}-1)}$
$=1+(\lceil q(\lfloor N_{m}/2\rfloor+1)/(q-1)\rceil-2)r$
$\geq 1+(q(\lfloor N_{m}/2\rfloor+1)/(q-1)-2)r$
$>1+\left(qN_{m}/(2(q-1))-2\right)r,$
which gives the desired result. ∎
The following lemma is a generalization of [32, Lemma 9] for determining some elements which are not coset leaders.
Lemma 2.8.
Assume $nr=(q^{m}-1)/s$. Then, $q^{i}+qt+j$ is not a coset leader modulo $nr$ for any integer $i$ with $(m+1)/2\leq i\leq\lfloor m-\log_{q}(s+1)\rfloor$, any integer $j$ with $1\leq j\leq q-1$ and any integer $t$ with
$0\leq t<\frac{q^{i}-j(q^{m-i}-1)-1}{q^{m-i+1}-q}.$
Moreover, all these $q^{i}+qt+j$ are pairwise distinct.
Proof.
First, we have
$(q^{i}+qt+j)q^{m-i}\equiv$ $(qt+j)q^{m-i}+1\pmod{q^{m}-1}$
$\equiv$ $(qt+j)q^{m-i}+1\pmod{nr}$.
In addition, we have
$(qt+j)q^{m-i}+1$
$=(qt+j)(q^{m-i}-1)+qt+j+1$
$<(\frac{q^{i}-j(q^{m-i}-1)-1}{q^{m-i}-1}+t)\cdot(q^{m-i}-1)+qt+j+1$
$=q^{i}-1+qt+j+1$
$=q^{i}+qt+j,$
and
$q^{i}+qt+j<q^{i}+\frac{q^{i}-j(q^{m-i}-1)-1}{q^{m-i}-1}+j=\frac{q^{m}-1}{q^{m-i}-1}\leq\frac{q^{m}-1}{s}=nr,$
where we need the condition $i\leq\lfloor m-\log_{q}(s+1)\rfloor$. Hence, $q^{i}+qt+j$ is not a coset leader modulo $nr$.
Due to the choice of $j$, if we fix $i$, then clearly $q^{i}+qt+j$ are pairwise distinct for various $j$ and $t$.
Finally, notice that for any $i$ with $(m+1)/2\leq i\leq\lfloor m-\log_{q}(s+1)\rfloor-1$, we have
$\begin{split}q^{i}+qt+j&<q^{i}+q\cdot\frac{q^{i}-(q^{2}-1)-1}{q^{3}-q}+q-1\\ &<2q^{i}\leq q^{i+1}.\end{split}$ (2.1)
Hence, we conclude that all these $q^{i}+qt+j$ are pairwise distinct. ∎
2.3 The case with two kinds of cyclotomic cosets
In this section, we assume that there are exactly two possible values of $l_{i}$, that is, $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integers $l,m$, where $l<m$ and $l\mid m$.
In this case, we can obtain better estimates about the coset leaders $\delta_{i}^{(l)}$ for some special $i$’s.
Lemma 2.9.
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$ and $N_{l}>1$. Then,
$\delta_{\lceil N_{l}/2\rceil+1}^{(l)}\geq\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
Proof.
First, since $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$, by Lemma 2.4 we have
$N_{l}=\frac{\gcd(n,(q^{l}-1)/r)}{l},$
and
$N_{m}=\frac{n-\gcd(n,(q^{l}-1)/r)}{m}.$
So, $N_{m}<n/m$, and then
$\frac{qN_{m}}{2(q-1)}\cdot r<\frac{q}{q-1}\cdot\frac{nr}{2m}.$
Hence, for our purpose it suffices to prove
$\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}\geq\frac{q}{q-1}\cdot\frac{nr}{2m}.$ (2.2)
Now, assume $N_{l}$ is even. Then, by Lemma 2.6 we have
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}>&(N_{l}/2+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\frac{\gcd(n,(q^{l}-1)/r)}{2l}\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\frac{nr}{2l}\geq\frac{nr}{m}\geq\frac{q}{q-1}\cdot\frac{nr}{2m},\end{split}$
where we also use $m\geq 2l$ (since $l\mid m$ and $l<m$). So, (2.2) holds in this case.
Finally, assume that $N_{l}$ is odd and $N_{l}>1$. Considering the case when $q=2$, we have $r=1$, which implies $n\in Z_{n,r}$ and the coset
$C_{n}^{(q,nr)}=\{n\}.$
So, in this case we have $l=1$ and $N_{l}=\gcd(n,1)=1$, which contradicts with $N_{l}>1$. Thus, we must have $q\geq 3$.
In addition, suppose $N_{l}=q=3$. Then, due to $N_{l}=\gcd(n,(q^{l}-1)/r)/l$, we have $N_{l}\mid n$. So, we have $q\mid n$, which contradicts with $\gcd(n,q)=1$. Hence, we must have either $N_{l}>3$ or $q>3$. Then, we get
$\frac{N_{l}}{N_{l}-1}\cdot\frac{q}{q-1}\leq\frac{3}{3-1}\cdot\frac{4}{4-1}=2.$ (2.3)
By Lemma 2.6, we obtain
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}>&(N_{l}/2-1/2+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\left(\frac{\gcd(n,(q^{l}-1)/r)}{2l}-\frac{1}{2}\right)\cdot\frac{nr^{2}}{\gcd(nr,q^{l}-1)}\\ =&\left(\frac{1}{2l}-\frac{r}{2\gcd(nr,q^{l}-1)}\right)nr\\ =&\left(\frac{1}{2l}-\frac{1}{2lN_{l}}\right)nr=\frac{1}{2l}\cdot\frac{N_{l}-1}{N_{l}}\cdot nr.\end{split}$
Then, combining this with (2.3) and $m\geq 2l$, we have
$\begin{split}\delta_{\lfloor N_{l}/2\rfloor+1}&>\frac{1}{2l}\cdot\frac{N_{l}-1}{N_{l}}\cdot nr\\ &\geq\frac{1}{m}\cdot\frac{q}{2(q-1)}\cdot nr=\frac{q}{q-1}\cdot\frac{nr}{2m}.\end{split}$
This gives (2.2), and then we complete the proof. ∎
In Lemma 2.9, if $N_{l}=1$, then the lower bound there for $\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}$ is not true in general; see Example 2.10.
Example 2.10.
Let $q=7$, $n=19$, $r=6$. Then, $m=3$, and the $7$-cyclotomic cosets in $Z_{19,6}$ are:
$C_{1}^{(7,114)}=\{1,7,49\},\quad C_{13}^{(7,114)}=\{13,91,67\},\quad C_{19}^{(7,114)}=\{19\},$
$C_{25}^{(7,114)}=\{25,61,85\},\quad C_{31}^{(7,114)}=\{31,103,37\},$
$C_{43}^{(7,114)}=\{43,73,55\},\quad C_{79}^{(7,114)}=\{79,97,109\}.$
So, $\{l_{i},i\in Z_{19,6}\}=\{1,3\}$, $N_{1}=1$ and $N_{3}=6$. Noticing that
$\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}=\delta_{1}^{(1)}=19$
and
$\frac{qN_{m}}{2(q-1)}\cdot r=\frac{7\times 6}{12}\times 6=21,$
we have
$\delta_{\lfloor N_{l}/2\rfloor+1}^{(l)}<\frac{qN_{m}}{2(q-1)}\cdot r.$
So, when $N_{l}=1$, we need to impose some extra condition to achieve such a lower bound.
Lemma 2.11.
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$, $N_{l}=1$ and $2m(q-1)\geq qlr$. Then,
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
Proof.
As in (2.2) and noticing $N_{l}=1$, for our purpose it suffices to prove
$\delta_{1}^{(l)}\geq\frac{q}{q-1}\cdot\frac{nr}{2m}.$ (2.4)
First, assume $q=2$, then we have $r=1$, which implies $n\in Z_{n,r}$ and the coset
$C_{n}^{(q,nr)}=\{n\}.$
So, we have $l=1$ and $N_{l}=\gcd(n,1)=1$. Hence,
$\delta_{1}^{(l)}=n>\frac{q}{q-1}\cdot\frac{nr}{2m},$
which implies (2.4) when $q=2$.
Now, assume $q>2$. By Lemma 2.4, we have $N_{l}=\gcd(n,(q^{l}-1)/r)/l$. Then, noticing $N_{l}=1$, we obtain
$\gcd(n,(q^{l}-1)/r)=l.$
Combining this with the condition $2m(q-1)\geq qlr$ and using Lemma 2.6, we have
$\begin{split}\delta_{1}^{(l)}\geq&\frac{nr}{\gcd(nr,q^{l}-1)}\\ =&\frac{n}{l}=\frac{2m}{l}\cdot\frac{n}{2m}\geq\frac{qlr}{l(q-1)}\cdot\frac{n}{2m}=\frac{q}{q-1}\cdot\frac{nr}{2m}.\end{split}$
Thus, (2.4) also holds in this case. This completes the proof. ∎
In Lemma 2.11, when $r=1,2\text{ or }3$, the extra condition
$2m(q-1)\geq qlr$
in fact always holds.
Corollary 2.12.
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume that $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$, $N_{l}=1$ and $r\leq 3$. Then,
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
Proof.
Since $r\mid q-1$, we have $r\leq q-1$. So,
$\frac{q}{q-1}\cdot r\leq\frac{r+1}{r}\cdot r=r+1.$ (2.5)
Noticing $m/l\geq 2$ and $r\leq 3$, we have
$2m/l\geq 4\geq r+1.$ (2.6)
Combining (2.5) with (2.6), we get
$2m/l\geq\frac{q}{q-1}\cdot r,$
that is,
$2m(q-1)\geq qlr.$
Thus, by Lemma 2.11, we have
$\delta_{1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r.$
When $l=1$ in Lemma 2.9, we can do a little better. One can see this by combining the following lemma with Lemma 2.7.
Lemma 2.13.
Write $nr=(q^{m}-1)/s$ for some positive integers $m,s$. Assume $\{l_{i}:i\in Z_{n,r}\}=\{1,m\}$ with $m\geq 2$. Then, if $N_{1}\geq 2$, we have
$\delta_{\lceil N_{1}/2\rceil+1}^{(1)}>\delta_{\lceil N_{m}/2\rceil}^{(m)};$
and if $N_{1}$ is even or $r\leq 2$, we have
$\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}>\delta_{\lfloor N_{m}/2\rfloor}^{(m)}.$
Proof.
Using Lemma 2.4 and noticing $\{l_{i}:i\in Z_{n,r}\}=\{1,m\}$, we have
$\begin{cases}N_{1}=\gcd(n,(q-1)/r),\\ N_{m}=(n-\gcd(n,(q-1)/r))/m.\end{cases}$ (2.7)
First, we estimate $\delta_{\lceil N_{1}/2\rceil+1}^{(1)}$. Combining (2.7) with Lemma 2.6, we have
$\begin{split}\delta_{\lceil N_{1}/2\rceil+1}^{(1)}\geq&(\lceil N_{1}/2\rceil+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+1\\ \geq&\frac{\gcd(n,(q-1)/r)}{2}\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+1\\ =&\frac{nr}{2}+1.\end{split}$ (2.8)
So, for our purpose we only need to prove
$\delta_{\lceil N_{m}/2\rceil}^{(m)}<\frac{nr}{2}+1.$ (2.9)
By contradiction, we suppose $\delta_{\lceil N_{m}/2\rceil}^{(m)}\geq\frac{nr}{2}+1$. Then, combining (2.7) with (2.8), we deduce that in the set $Z_{n,r}$, the total number of elements in all the $q$-cyclotomic cosets with coset leader less than $\frac{nr}{2}+1$ is at most
$\begin{split}&\lceil N_{1}/2\rceil+m\left(\lceil N_{m}/2\rceil-1\right)\\ &\leq\frac{\gcd(n,(q-1)/r)}{2}+\frac{1}{2}+m\cdot\frac{n-\gcd(n,(q-1)/r)}{2m}-\frac{m}{2}\\ &=\frac{n}{2}-\frac{m-1}{2}.\end{split}$ (2.10)
However, in the set $Z_{n,r}$ the number of elements less than $\frac{nr}{2}+1$ is at least
$\lceil n/2\rceil\geq n/2,$
which contradicts with (2.10) (noticing $m\geq 2$). Hence, (2.9) is true, and thus we get
$\delta_{\lceil N_{1}/2\rceil+1}^{(1)}>\delta_{\lceil N_{m}/2\rceil}^{(m)}.$
Now, it remains to consider the case when $N_{1}$ is odd and $r\leq 2$. Since $N_{1}$ is odd, by Lemma 2.6 we obtain
$\begin{split}&\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}\\ \geq&(\lfloor N_{1}/2\rfloor+1-1)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+\frac{nr}{\gcd(nr,q-1)}\\ =&\left(\frac{\gcd(n,(q-1)/r)}{2}-\frac{1}{2}\right)\cdot\frac{nr^{2}}{\gcd(nr,q-1)}+\frac{nr}{\gcd(nr,q-1)}\\ =&\frac{nr}{2}+\frac{nr(2-r)}{2\gcd(nr,q-1)}\geq\frac{nr}{2},\end{split}$ (2.11)
where the last inequality follows from the assumption $r\leq 2$. Hence, when $N_{1}$ is odd and $r\leq 2$, for our purpose we only need to prove
$\delta_{\lfloor N_{m}/2\rfloor}^{(m)}<\frac{nr}{2}$ (2.12)
By contradiction, suppose $\delta_{\lfloor N_{m}/2\rfloor}^{(m)}\geq nr/2$. Then, combining (2.7) with (2.11), we deduce that in the set $Z_{n,r}$, the total number of elements in the $q$-cyclotomic cosets with coset leader less than $nr/2$ is at most
$\begin{split}&\lfloor N_{1}/2\rfloor+m\left(\lfloor N_{m}/2\rfloor-1\right)\\ &\leq\frac{\gcd(n,(q-1)/r)}{2}+m\cdot\frac{n-\gcd(n,(q-1)/r)}{2m}-m\\ &=\frac{n}{2}-m.\end{split}$ (2.13)
However, in the set $Z_{n,r}$ the number of elements less than $nr/2$ is at least
$\lfloor n/2\rfloor-1\geq\frac{n}{2}-\frac{3}{2},$
which contradicts with (2.13) (noticing $m\geq 2$). Therefore, (2.12) is true, and so, we obtain
$\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}>\delta_{\lfloor N_{m}/2\rfloor}^{(m)}.$
This completes the proof. ∎
We remark that if $N_{1}$ is odd and $r>2$, then $\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}$ may be less than $\delta_{\lfloor N_{m}/2\rfloor}^{(m)}$; see Example 2.14.
Example 2.14.
Choosing $q=19$, $n=127$ and $r=18$, we have $nr=(q^{m}-1)/s$ with $m=3$ and $s=3$, $\{l_{i},i\in Z_{127,18}\}=\{1,3\}$, $N_{1}=1$ and $N_{3}=42$. Moreover, $\delta_{1}^{(1)}=127$ and $\delta_{21}^{(3)}=451$. So,
$\delta_{\lfloor N_{1}/2\rfloor+1}^{(1)}=\delta_{1}^{(1)}<\delta_{21}^{(3)}.$
2.4 Cyclotomic values
Recall that for any positive integer $m$, the $m$-th cyclotomic polynomial is defined to be
$\Phi_{m}(x)=\prod_{\begin{subarray}{c}j=1\\ \gcd(j,m)=1\end{subarray}}^{m}(x-\zeta_{m}^{j})\in\mathbb{Z}[x],$
where $\zeta_{m}$ is a primitive $m$-th root of unity. Besides, $\Phi_{m}(x)$ can be expressed as
$\Phi_{m}(x)=\prod_{d\mid m}(x^{d}-1)^{\mu(m/d)},$ (2.14)
where $\mu$ is the Möbius function.
The following result is in [37, Proposition 2.8] about $\Phi_{m}(1)$.
Lemma 2.15 ([37]).
If $m$ has at least two distinct prime divisors, then $\Phi_{m}(1)=1$.
Next, we determine $\gcd(\Phi_{m}(q),q-1)$.
Lemma 2.16.
We have
$\gcd(\Phi_{m}(q),q-1)$
[…]
Proof.
First, we assume that $m$ has at least two distinct prime factors. By contradiction, suppose $\gcd(\Phi_{m}(q),q-1)>1$. Then, there exists a prime $p$ such that $p\mid\gcd(\Phi_{m}(q),q-1)$. However, by Lemma 2.15 and noticing $p\mid q-1$, we have
$\Phi_{m}(q)\equiv\Phi_{m}(1)\equiv 1\pmod{p}$,
which contradicts with $p\mid\Phi_{m}(q)$. Thus, $\gcd(\Phi_{m}(q),q-1)=1$. This proves the first part of the lemma.
Now, we assume $m=p^{j+1}$ for some non-negative integer $j$. By (2.14), we obtain
$\Phi_{m}(q)$ $=\frac{q^{p^{j+1}}-1}{q^{p^{j}}-1}$
$=q^{p^{j}(p-1)}+q^{p^{j}(p-2)}+\cdots+q^{p^{j}}+1$
$\equiv 1+1+\cdots+1+1\equiv p\pmod{q-1}$.
So, we have
$\gcd(\Phi_{m}(q),q-1)=\gcd(q-1,p).$
This proves the second part of the lemma. ∎
2.5 The BCH bound
The BCH lower bound is an important bound for the minimun distance of BCH codes, which can be extended for constacyclic codes; see [18, Theorem 2.2].
Lemma 2.17 ([18]).
If the defining set $T$ of a $\lambda$-constacyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ includes $1,1+r,1+2r,\dots,1+dr$, then the minimun distance $d(\mathcal{C})\geq d+2$.
3 Some general constructions for constacyclic codes
In this section, motivated by [32] and generalizing the constructions in [8, 20, 23, 31, 32], we propose a general construction for $q$-ary constacyclic codes. First we classify the $q$-cyclotomic cosets in $Z_{n,r}$ according to their sizes (that is, two cosets are in the same class if and only if they have the same size), and then we take the defining set $T$ as the union of the first half of cosets in each class.
Recall that $q$ is a prime power, $n$ is a positive integer with $\gcd(n,q)=1$, and $r\mid q-1$.
For simplicity, we still denote $N_{l}=N_{l}^{(q,nr)}$ and $\delta_{i}^{(l)}=\delta_{i}^{(q,nr,l)}$ for any positive integer $l$. By definition, $N_{l}$ is the number of $q$-cyclotomic cosets with $l$ elements in the set $Z_{n,r}$, and $\delta_{i}^{(l)}$ is the $i$-th coset leader (in ascending order) in those $q$-cyclotomic cosets with $l$ elements in $Z_{n,r}$.
Recall that for any $i\in Z_{n,r}$, $l_{i}=|C_{i}^{(q,nr)}|$, which is the size of the $q$-cyclotomic coset generated by $i$ modulo $nr$.
Theorem 3.1.
Write $nr=(q^{m}-1)/s$, and assume $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ for some positive integer $l<m$. Then, an infinite family of $q$-ary constacyclic $n,n-l\lceil N_{l}/2\rceil-m\lceil N_{m}/2\rceil,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed, where
$N_{l}=\gcd(n,(q^{l}-1)/r)/l\quad\textrm{and}\quad N_{m}=(n-\gcd(n,(q^{l}-1)/r))/m.$
Besides, if
  • (1)
    $N_{l}>1$, or
  • (2)
    $N_{l}=1$ and $2m(q-1)\geq qlr$, or
  • (3)
    $r\leq 3$,
then an infinite family of constacyclic $n,n-l\lfloor N_{l}/2\rfloor-m\lfloor N_{m}/2\rfloor,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed.
Proof.
Since $\{l_{i}:i\in Z_{n,r}\}=\{l,m\}$ with $l<m$, by Lemma 2.4 we have
$N_{l}=\gcd(n,(q^{l}-1)/r)/l\quad\textrm{and}\quad N_{m}=(n-\gcd(n,(q^{l}-1)/r))/m.$
Take the defining set $T$ as the union of the first $\lceil N_{l}/2\rceil$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $l$ elements and the first $\lceil N_{m}/2\rceil$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements. So, there are $l\lceil N_{l}/2\rceil+m\lceil N_{m}/2\rceil$ integers in $T$, which implies that the dimension of the corresponding constacyclic code $\mathcal{C}$ is
$n-l\lceil N_{l}/2\rceil-m\lceil N_{m}/2\rceil.$
Next, we prove the desired lower bound for the minimun distance $d(\mathcal{C})$. Since $d(\mathcal{C})\geq 2$, we can assume $N_{m}>4$ without loss of generality. When $N_{l}>1$, by Lemma 2.7 and Lemma 2.9, we have
$\delta_{\lceil N_{l}/2\rceil+1}^{(l)}>\frac{qN_{m}}{2(q-1)}\cdot r>1+\left(\frac{qN_{m}}{2(q-1)}-2\right)r$
and
$\delta_{\lceil N_{m}/2\rceil+1}^{(m)}>1+\left(\frac{qN_{m}}{2(q-1)}-2\right)r.$
By construction, this means that for any non-negative integer $i\leq\frac{qN_{m}}{2(q-1)}-2$, $1+ir\in T$. So, by Lemma 2.17 we get
$d(\mathcal{C})\geq\left\lfloor\frac{qN_{m}}{2(q-1)}\right\rfloor.$
Similarly, the case when $N_{l}=1$ follows from Lemma 2.9 and Lemma 2.17. So, we prove the first part of the theorem.
Now, take the defining set $T$ as the union of the first $\lfloor N_{l}/2\rfloor$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $l$ elements and the first $\lfloor N_{m}/2\rfloor$ $q$-cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements. Then, under one of the conditions (1), (2) and (3) and using Lemmas 2.9 and 2.11 and Corollary 2.12, an infinite family of constacyclic $n,n-l\lfloor N_{l}/2\rfloor-m\lfloor N_{m}/2\rfloor,\geq\lfloor\frac{qN_{m}}{2(q-1)}\rfloor]$ codes is constructed. ∎
We remark that the quantity $N_{m}$ in Theorem 3.1 is greater than $cn/\log_{q}n$ for some constant $c$ which does not depend on $m$. This means that the minimum distance $d(\mathcal{C})$ is also greater than $cn/\log_{q}n$.
The following example suggests that when $r=1$, the constructed codes in Theorem 3.1 may be not always BCH codes.
Example 3.2.
Let $q=5$, $n=26$, and $r=1$. Then, there are eight $5$-cyclotomic cosets modulo $26$:
$C_{1}^{(5,26)}=\{1,5,25,21\},\quad C_{2}^{(5,26)}=\{2,10,24,16\},$
$C_{3}^{(5,26)}=\{3,15,23,11\},\quad C_{4}^{(5,26)}=\{4,20,22,6\},$
$C_{7}^{(5,26)}=\{7,9,19,17\},\quad C_{8}^{(5,26)}=\{8,14,18,12\},$
$C_{13}^{(5,26)}=\{13\},\qquad\qquad C_{26}^{(5,26)}=\{26\}.$
Then, $N_{1}=2,N_{4}=6$. So, in Theorem 3.1, the defining set $T$ is
$C_{1}^{(5,26)}\cup C_{2}^{(5,26)}\cup C_{3}^{(5,26)}\cup C_{13}^{(5,26)},$
and the corresponding code is not a BCH code.
Now, we want to establish a better lower bound for minimum distance in some special cases.
Theorem 3.3.
Write $nr=(q^{m}-1)/s$, and assume $\{l_{i}:i\in Z_{n,r}\}=\{1,m\}$ for some positive integer $m>1$. Then, for the first family of $q$-ary constacyclic codes in Theorem 3.1, the lower bound on their minimum distances can be
$\left\lceil\frac{q\lceil N_{m}/2\rceil}{q-1}\right\rceil+\Lambda_{m},$
where $N_{m}$ is given in Theorem 3.1,
$\Lambda_{m}=\left\lceil\frac{q|S_{m}|}{q-1}\right\rceil-1,$
and
$\begin{split}S_{m}=\Big\{(i,j,t):\ &\frac{m+1}{2}\leq i\leq\lfloor m-\log_{q}(s+1)\rfloor,\\ &1\leq j\leq q-1,\ 0\leq t<\frac{q^{i}-j(q^{m-i}-1)-1}{q^{m-i+1}-q},\\ &q^{i}+qt+j\equiv 1\pmod{r},\\ &q^{i}+qt+j\leq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2\right)r\Big\};\end{split}$
and if moreover $r\leq 2$, then for the second family of constacyclic codes in Theorem 3.1, the lower bound on their minimum distances can be
$\left\lceil\frac{q\lfloor N_{m}/2\rfloor}{q-1}\right\rceil+\Lambda_{m}^{\prime},$
where
$\Lambda_{m}^{\prime}=\left\lceil\frac{q|S_{m}^{\prime}|}{q-1}\right\rceil-1,$
and the set $S_{m}^{\prime}$ is obtained from $S_{m}$ by replacing $\lceil N_{m}/2\rceil$ with $\lfloor N_{m}/2\rfloor$.
Proof.
Recall that for the first family of constacyclic codes in Theorem 3.1, their defining set $T$ is the union of the first $\lceil N_{1}/2\rceil$ cyclotomic cosets contained in $Z_{n,r}$ with one element and the first $\lceil N_{m}/2\rceil$ cyclotomic cosets contained in $Z_{n,r}$ with $m$ elements.
By construction and by the first part of Lemma 2.13, we know that when $N_{1}\geq 2$, all the cyclotomic cosets of $Z_{n,r}$ with coset leader at most $\delta_{\lceil N_{m}/2\rceil}^{(m)}$ are in $T$. When $N_{1}=1$, clearly this fact also holds.
Moreover, by Lemma 2.6 we get
$\delta^{(m)}_{\lceil N_{m}/2\rceil}$ $\geq\frac{nr(1+(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2)r)}{\gcd(nr,q^{m}-1)}$
$=1+(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2)r.$
Notice that the above lower bound is achieved only when all the elements not divisible by $q$ and at most $1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2\right)r$ in $Z_{n,r}$ are coset leaders (because in proving Lemma 2.6, we apply Lemma 2.5). However, this does not always hold. By Lemma 2.8, we know that there are at least $|S_{m}|$ such elements which are not coset leaders. Hence, combining this with Lemma 2.5, we obtain
$\begin{split}\delta^{(m)}_{\lceil N_{m}/2\rceil}&\geq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-1\right)r+(\lceil q|S_{m}|/(q-1)\rceil-2)r\\ &\geq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil+\lceil q|S_{m}|/(q-1)\rceil-3\right)r,\end{split}$ (3.1)
where the set $S_{m}$ has been defined in the theorem.
Hence, using (3.1) and applying Lemma 2.17, we get the desired lower bound for the minimum distances of the first family of constacyclic codes in Theorem 3.1.
Finally, when $r\leq 2$, using the second part of Lemma 2.13 and applying similar arguments as the above, we get the desired lower bound for the minimum distances of the second family of constacyclic codes in Theorem 3.1. ∎
We remark that the sizes of the sets $S_{m}$ and $S_{m}^{\prime}$ in Theorem 3.3 can be easily computed by computer. We give an estimate for $|S_{m}|$ below (the method also works for estimating $|S_{m}^{\prime}|$), although it is not always good according to numerical data. Let
$e(m)=\min\{\lfloor m-\log_{q}(s+1)\rfloor,\ \lfloor\log_{q}(1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2\right)r)\rfloor\}.$
For any $i$ with $(m+1)/2\leq i\leq e(m)-1$ and for any $j,t$ described in $S_{m}$, by (2.1) and noticing the choice of $e(m)$, we have
$q^{i}+qt+j<q^{i+1}\leq q^{e(m)}\leq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2\right)r.$
Therefore, using the second part of Lemma 2.8, we obtain
$|S_{m}|\geq\sum_{i=(m+1)/2}^{e(m)-1}\sum_{t=1}^{q-1}\left\lfloor\frac{q^{i}-t(q^{m-i}-1)-1}{r(q^{m-i+1}-q)}\right\rfloor,$
where $r$ appears in the denomenator due to the condition $q^{i}+qt+j\equiv 1\pmod{r}$.
The data in Table 3.1 suggest that the improvement of the lower bound in Theorem 3.3 is indeed meaningful (compared to Theorem 3.1).
Table 3.1: The lower bound in Theorem 3.3
$q$ $m$ $s$ $r$ $\left\lfloor\frac{qN_{m}}{2(q-1)}\right\rfloor$ $\left\lceil\frac{q\lceil N_{m}/2\rceil}{q-1}\right\rceil+\Lambda_{m}$ Bose distance
3 5 1 1 36 38 41
3 11 1 1 12078 12359 14530
3 11 1 2 6039 6179 7266
3 13 1 2 45990 47263 55357
4 7 1 1 1560 1631 1835
4 7 1 3 520 543 613
5 7 2 1 3487 3622 4157
5 7 2 2 1743 1811 2080
7 7 3 1 22876 23232 27299
7 7 3 2 11438 11616 13651
4 Constacyclic codes with prime length
Recall that $q$ is a prime power, $n$ is a positive integer with $\gcd(n,q)=1$ and $r\mid q-1$. In this section, $n$ is a prime.
Theorem 4.1.
Assume that $n$ is a prime with $n\nmid q-1$, and let $m=\mathrm{ord}_{n}(q)$. Then, an infinite family of constacyclic $n,n-1-m\lceil\frac{n-1}{2m}\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2m)\rceil}{q-1}\right\rceil+\Lambda_{m}]$ codes is constructed. Besides, if $2m(q-1)\geq qr$, an infinite family of constacyclic $n,n-m\lfloor\frac{n-1}{2m}\rfloor,\geq\lfloor\frac{q(n-1)}{2m(q-1)}\rfloor]$ codes is constructed.
Proof.
Since $n$ is a prime and $n\nmid q-1$, we have $\gcd(n,q-1)=1$, and so $\gcd(n,r)=1$ (noticing $r\mid q-1$), which implies
$\mathrm{ord}_{nr}(q)=\mathrm{ord}_{n}(q)\mathrm{ord}_{r}(q)=\mathrm{ord}_{n}(q).$
Let $m=\mathrm{ord}_{n}(q)$. Then, combining Lemma 2.1 with Lemma 2.4 and noticing that $n$ is a prime, we obtain that
$\{l_{i}:i\in Z_{n,r}\}=\{1,m\},$
and
$N_{1}=\gcd(n,(q-1)/r)=1\qquad\textrm{and}\qquad N_{m}=\frac{n-1}{m}.$
Hence, the desired results follow directly from Theorems 3.1 and $3.$ ∎
We remark that in Theorem 4.1, if furthermore $r\leq 2$, then by Theorem 3.3, we know that for the second family of constacyclic codes, the lower bound $\lfloor\frac{q(n-1)}{2m(q-1)}\rfloor$ can be replaced by
$\left\lceil\frac{q\lfloor(n-1)/(2m)\rfloor}{q-1}\right\rceil+\Lambda_{m}^{\prime},$
where $\Lambda_{m}^{\prime}$ has been defined in Theorem 3.3.
We also remark that in Theorem 4.1, when (n1)/m(n-1)/m is even, the two families of constacyclic codes have dimensions (n1)/2(n-1)/2 and $(n+1)/2$ respectively.
The lower bounds on minimum distances in Theorem 4.1 depend on the quotient $\frac{n-1}{\mathrm{ord}_{n}(q)}$ when $q$ is fixed. This quotient is from 1 to n1n-1. Particularly, when $q$ is a prime, it is conjectured that there are infinitely many primes $p$ such that $\frac{q^{p}-1}{q-1}$ is also a prime (called the generalized repunit prime). For this case, we put $n=\frac{q^{p}-1}{q-1}$, and then $\mathrm{ord}_{n}(q)=p$, which implies that the lower bounds are of the form $cn/\log_{q}n$ for some positive constant $c$.
In Table LABEL:tab:nprimec, we present some codes constructed in Theorem 4.1 with prime length $n$ and having optimal or best-known parameters. In this table, $m=\mathrm{ord}_{n}(q)$. We emphasize that those codes marked with “*” in the parameters are constructed by using the ceiling function (that is, from the first family of constacyclic codes in the theorem), and the others are constructed by using the floor function (that is, from the second family of constacyclic codes in the theorem). In addition, the column “Lower bound” records the best lower bound in this paper for the minimum distance of the relevant code (here, actually from Theorems 3.3 and 4.1). We follow these rules in all the other tables of constacyclic codes in this paper.
Table 4.1: Codes in Theorem 4.1
$q$ $m$ $n$ $r$ Lower Bose Parameter Optimality
bound distance
22 2020 $41$ 11 22 44 $41,20,10]^{*}$ best-known
22 2020 $41$ 11 22 44 $[41,21,9]$ best-known
22 $23$ $47$ 11 22 55 $47,23,12]^{*}$ best-known
22 $23$ $47$ 11 22 55 $[47,24,11]$ best-known
33 33 $13$ 11 33 44 $13,6,6]^{*}$ optimal
33 33 $13$ 22 33 44 $13,6,6]^{*}$ optimal
33 $12$ $73$ 11 44 55 $73,36,18]^{*}$ best-known
33 $12$ $73$ 11 44 55 $[73,37,17]$ best-known
33 $12$ $73$ 22 55 55 $73,36,18]^{*}$ best-known
33 $12$ $73$ 22 55 77 $[73,37,17]$ best-known
33 $23$ $47$ 11 22 66 $47,23,15]^{*}$ best-known
33 $23$ $47$ 11 22 55 $[47,24,14]$ best-known
33 $23$ $47$ 22 22 66 $47,23,15]^{*}$ best-known
33 $23$ $47$ 22 22 55 $[47,24,14]$ best-known
55 33 $31$ 11 77 88 $[31,16,10]$ optimal
55 33 $31$ 22 77 88 $[31,16,10]$ optimal
55 33 $31$ 44 77 88 $[31,16,10]$ optimal
55 $14$ $29$ 11 22 55 $29,14,12]^{*}$ optimal
55 $14$ $29$ 11 22 55 $[29,15,11]$ optimal
55 $14$ $29$ 22 22 44 $29,14,12]^{*}$ optimal
55 $14$ $29$ 22 22 44 $[29,15,11]$ optimal
55 $14$ $29$ 44 22 55 $29,14,12]^{*}$ optimal
55 $14$ $29$ 44 22 55 $[29,15,11]$ optimal
5 Constacyclic codes with $n=\frac{q^{p}-1}{rs}$
In this section, we assume $nr=\frac{q^{p}-1}{s}$ for some prime $p$ and some positive integer $s$. Recall that $r\mid q-1$.
Theorem 5.1.
Assume that $nr=(q^{p}-1)/s$ for some prime $p$, and $\gcd(r,nr/\gcd(nr,q-1))=1$. Then, an infinite family of constacyclic $n,n-\lceil N_{1}/2\rceil-p\lceil N_{p}/2\rceil,\geq\left\lceil\frac{q\lceil N_{p}/2\rceil}{q-1}\right\rceil+\Lambda_{p}]$ codes is constructed, where
$N_{1}=\gcd(n,(q-1)/r),\qquad N_{p}=(n-\gcd(n,(q-1)/r))/p,$
and $\Lambda_{p}$ has been defined in Theorem 3.3. Moreover, if
  • (1)
    $N_{1}>1$, or
  • (2)
    $N_{1}=1$ and $2p(q-1)\geq qr$, or
  • (3)
    $r\leq 3$,
then an infinite family of constacyclic $n,n-\lfloor N_{1}/2\rfloor-p\lfloor N_{p}/2\rfloor,\geq\lfloor\frac{qN_{p}}{2(q-1)}\rfloor]$ codes is constructed.
Proof.
Since $nr=(q^{p}-1)/s$ and $p$ is a prime, by Lemma 2.1 we have that for any $i\in Z_{n,r}$, $l_{i}=1$ or $p$, where $l_{i}=|C_{i}^{(q,nr)}|$. Combining this with Lemma 2.4 (noticing $\gcd(r,nr/\gcd(nr,q-1))=1$), we have
$N_{l}=\begin{cases}\gcd(n,(q-1)/r),&l=1,\\ \frac{n-\gcd(n,(q-1)/r)}{p},&l=p.\end{cases}$
Then, the desired results follow directly from Theorems 3.1 and $3.$ ∎
We remark that in Theorem 5.1, if furthermore $r\leq 2$, then by Theorem 3.3, we know that for the second family of constacyclic codes, the lower bound $\lfloor\frac{qN_{p}}{2(q-1)}\rfloor$ can be replaced by
$\left\lceil\frac{q\lfloor N_{p}/2\rfloor}{q-1}\right\rceil+\Lambda_{p}^{\prime},$
where $\Lambda_{p}^{\prime}$ has been defined in Theorem 3.3.
In Theorem 5.1, there are some special cases where such constacyclic codes have dimension (n1)/2(n-1)/2, or $n/2$, or $(n+1)/2$. We state some of them as examples.
Example 5.2.
In Theorem 5.1, if moreover $n$ is even, $p$ is an odd prime and $2r\mid q-1$, then both $N_{1}$ and $N_{p}$ are even, and so the first family of constacyclic codes has dimension $n/2$.
Example 5.3.
In Theorem 5.1, if moreover $n$ is odd and $p$ is an odd prime, then $N_{1}$ is odd and $N_{p}$ is even, and so, the first family of constacyclic codes has dimension
$n-\frac{N_{1}+1}{2}-\frac{pN_{p}}{2}=\frac{n-1}{2},$
and the second family of constacyclic codes has dimension
$n-\frac{N_{1}-1}{2}-\frac{pN_{p}}{2}=\frac{n+1}{2}.$
In Table LABEL:tab:qoddprceilc, we list some good codes constructed in Theorem 5.1. In the table, $d_{\text{best}}$ stands for the maximal minimum distance of all known $q$-ary linear codes with relevant length and dimension, and its value is from [15].
Table 5.1: Codes in Theorem 5.1
$q$ $p$ $s$ $r$ Lower Bose Parameter Optimality
bound distance
2 11 89 1 2 6 $23,11,8]^{*}$ optimal
2 11 89 1 2 5 $[23,12,7]$ optimal
3 3 2 1 3 5 $13,6,6]^{*}$ optimal
3 3 2 1 3 4 $[13,7,4]$ almost-optimal
3 3 1 1 6 6 $26,13,8]^{*}$ best-known
3 5 22 1 2 5 $11,5,6]^{*}$ optimal
3 5 22 1 1 4 $[11,6,5]$ optimal
3 5 11 2 2 4 $11,5,6]^{*}$ optimal
3 5 11 2 1 4 $[11,6,5]$ optimal
3 5 11 1 3 4 $22,11,7]^{*}$ almost-optimal
3 11 7702 1 2 6 $23,11,9]^{*}$ optimal
3 11 7702 1 1 5 $[23,12,8]$ optimal
3 11 3851 2 1 6 $23,11,9]^{*}$ optimal
3 11 3851 1 1 5 $[23,12,8]$ optimal
4 3 3 1 4 6 $21,10,8]^{*}$ best-known
4 3 1 1 13 14 $[63,32,15]$ $d_{\text{best}}=16$
5 3 1 1 25 32 $124,62,32]^{*}$ $d_{\text{best}}=33$
5 5 142 1 3 5 $22,11,8]^{*}$ best-known
5 5 71 2 3 5 $22,11,8]^{*}$ best-known
7 3 18 1 4 6 $19,9,8]^{*}$ almost-optimal
7 3 6 3 3 6 $[19,10,7]$ almost-optimal
7 7 9466 3 3 4 $29,14,12]^{*}$ best-known
7 7 9466 3 2 4 $[29,15,11]$ best-known
6 Constacyclic codes with $n=\Phi_{p^{b}}(q)$
Recall that $q$ is a prime power, and $r\mid q-1$. Let $p$ be a prime and $b$ be a positive integer. In this section, we construct some infinite families of constacyclic codes with length
$n=\Phi_{p^{b}}(q)=\frac{q^{p^{b}}-1}{q^{p^{b-1}}-1}.$
Later on, we will see that in this case the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$, always has exactly two elements. So, we can use Theorem 3.1 to construct constacyclic codes having good lower bounds for their minimum distances.
First, we consider the case when $p\nmid q-1$.
Theorem 6.1.
Assume $p\nmid q-1$ and $n=(q^{p^{b}}-1)/(q^{p^{b-1}}-1)$. Then, an infinite family of constacyclic $n,n-1-p^{b}\lceil(n-1)/(2p^{b})\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p^{b}(q-1)\geq qr$, an infinite family of constacyclic $n,n-p^{b}\lfloor(n-1)/(2p^{b})\rfloor,\geq\lfloor\frac{q(n-1)}{2p^{b}(q-1)}\rfloor]$ codes is constructed.
Proof.
First, we want to determine $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. Since $r\mid q-1$ and $b\geq 1$, we write
$nr=\frac{q^{p^{b}}-1}{(q^{p^{b-1}}-1)/r},$ (6.1)
where the denomenator is a positive integer. Combining (6.1) with Lemma 2.1, we have
$l_{i}\mid p^{b}\quad\textrm{for}\quad i\in Z_{n,r}.$
For $j=0,1,\ldots,b-1$, we have
[…]
which implies
$\gcd(n,q^{p^{j}}-1)=\gcd(p,q^{p^{j}}-1).$
If $p\mid q$, clearly we have $\gcd(p,q^{p^{j}}-1)=1$. In addition, if $p\nmid q$, suppose $\gcd(p,q^{p^{j}}-1)=p$, then we have $p\mid q-1$ (because $p\mid q^{p^{j}}-1$ and $p\mid q^{p-1}-1$), which contradicts with the assumption $p\nmid q-1$, and thus, we must have $\gcd(p,q^{p^{j}}-1)=1$. Therefore, we always have
$\gcd(n,q^{p^{j}}-1)=\gcd(p,q^{p^{j}}-1)=1\quad\textrm{for}\quad j=0,1,\ldots,b-1.$
Hence, we get
$\gcd(n,q^{p^{j}}-1)=1\qquad\textrm{for}\qquad j=0,1,\ldots,b-1,$ (6.3)
which also implies (noticing $r\mid q-1$)
$\gcd\left(r,\frac{nr}{\gcd(nr,q^{j}-1)}\right)=\gcd\left(r,\frac{n}{\gcd(n,(q^{j}-1)/r)}\right)=1.$
Thus, applying Lemma 2.3 and (6.3) and using the definition of the Möbius function, we obtain
$N_{1}$ $=\gcd(n,(q-1)/r)=1,$
$N_{p^{b}}$ $=\Big(\sum_{j=0}^{b}\mu(p^{b-j})\gcd(n,(q^{p^{j}}-1)/r)\Big)/p^{b}$
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}$
$=\frac{n-1}{p^{b}}.$
Moreover, noticing $N_{1}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have
$\{l_{i}:\,i\in Z_{n,r}\}=\{1,p^{b}\}.$
Then, the desired result follows directly from Theorems 3.1 and $3.$ ∎
Now, we consider the case when $p\mid\frac{q-1}{r}$.
Theorem 6.2.
Assume $p\mid(q-1)/r$ and $n=(q^{p^{b}}-1)/(q^{p^{b-1}}-1)$. Then, an infinite family of constacyclic $n,n-\lceil p/2\rceil-p^{b}\lceil\frac{n-p}{2p^{b}}\rceil,\geq\left\lceil\frac{q\lceil(n-p)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p^{b}(q-1)\geq qr$, an infinite family of constacyclic $n,n-\lfloor p/2\rfloor-p^{b}\lfloor\frac{n-p}{2p^{b}}\rfloor,\geq\lfloor\frac{q(n-p)}{2p^{b}(q-1)}\rfloor]$ codes is constructed.
Proof.
First, we want to determine the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. As before, we have $l_{i}\mid p^{b}$ for any $i\in Z_{n,r}$.
Combining (6.2) with the assumption $p\mid(q-1)/r$, we have that for $j=0,1,\ldots,b-1$,
$\gcd(n,(q^{p^{j}}-1)/r)=\gcd(p,(q^{p^{j}}-1)/r)=p.$ (6.4)
In addition, by (6.2) and noticing $r\mid q-1$, we have
$\gcd(n,r)=\gcd(p,r).$
So, $\gcd(n,r)=1$ or $p$. If $\gcd(n,r)=1$, clearly we have $\gcd(n/p,r)=1$. If $\gcd(n,r)=p$, then noticing $p\mid(q-1)/r$, we can write $q=ap^{2}+1$ for some positive integer $a$, and so,
[…]
which implies $p\nmid n/p$. Thus, when $\gcd(n,r)=p$, we have $p\nmid n/p$, and so $\gcd(n/p,r)=1$. Hence, we always have
$\gcd(n/p,r)=1.$ (6.5)
Therefore, combining (6.4) with (6.5), we get that for $j=0,1,\ldots,b-1$,
$\begin{split}\gcd(r,\frac{nr}{\gcd(nr,q^{p^{j}}-1)})&=\gcd(r,\frac{n}{\gcd(n,(q^{p^{j}}-1)/r)})\\ &=\gcd(r,n/p)=1.\end{split}$ (6.6)
Now, applying Lemma 2.3 with (6.4) and (6.6) and using the definition of the Möbius function, we obtain
$N_{1}$ $=\gcd(n,(q-1)/r)=p,$
$N_{p^{b}}$ $=\Big(\sum_{j=0}^{b}\mu(p^{b-j})\gcd(n,(q^{p^{j}}-1)/r)\Big)/p^{b}$
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}$
$=\frac{n-p}{p^{b}}.$
Moreover, noticing $N_{1}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have
$\{l_{i}:\,i\in Z_{n,r}\}=\{1,p^{b}\}.$
Then, the desired result follows directly from Theorems 3.1 and $3.$ ∎
Finally, we consider the case when $p\mid q-1$ and $p\nmid(q-1)/r$.
Theorem 6.3.
Assume $p\mid q-1$, $p\nmid(q-1)/r$, $b\geq 2$ and $n=(q^{p^{b}}-1)/(q^{p^{b-1}}-1)$. Then, an infinite family of constacyclic $n,n-p-p^{b}\lceil\frac{n-p}{2p^{b}}\rceil,\geq\left\lceil\frac{q\lceil(n-p)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p^{b-1}(q-1)\geq qr$, an infinite family of constacyclic $n,n-p^{b}\lfloor\frac{n-p}{2p^{b}}\rfloor,\geq\lfloor\frac{q(n-p)}{2p^{b}(q-1)}\rfloor]$ codes is constructed.
Proof.
First, we want to determine the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. As before, we have $l_{i}\mid p^{b}$ for any $i\in Z_{n,r}$.
Since $p\mid q-1$ and $p\nmid(q-1)/r$, we must have $p\mid r$.
Combining (6.2) with the assumption $p\nmid(q-1)/r$, we have
$\gcd(n,(q-1)/r)=\gcd(p,(q-1)/r)=1.$ (6.7)
Similarly, combining (6.2) with the assumption $p\mid q-1$, for $j=1,\ldots,b-1$ we have
$\begin{split}\gcd(n,(q^{p^{j}}-1)/r)&=\gcd(p,(q^{p^{j}}-1)/r)\\ &=\gcd(p,\frac{q-1}{r}(q^{p^{j}-1}+\cdots+q+1))\\ &=\gcd(p,q^{p^{j}-1}+\cdots+q+1)=p.\end{split}$ (6.8)
In addition, since $p\mid q-1$, we write $q=ap+1$. Then,
[…]
where the last congruence follows from the assumption $b\geq 2$. So, we have $p\nmid n/p$. Notice that $\gcd(n,r)=\gcd(p,r)=1$ or $p$. Then, we must have
$\gcd(n/p,r)=1.$ (6.9)
Then, using (6.7) and noticing $p\mid r$, we obtain
$\begin{split}\gcd\left(r,\frac{nr}{\gcd(nr,q-1)}\right)&=\gcd\left(r,\frac{n}{\gcd(n,(q-1)/r)}\right)\\ &=\gcd(r,n)=\gcd(r,p)=p.\end{split}$ (6.10)
For $j=1,\ldots,b-1$, using (6.8) and (6.9), we have
$\begin{split}\gcd\left(r,\frac{nr}{\gcd(nr,q^{p^{j}}-1)}\right)&=\gcd\left(r,\frac{n}{\gcd(n,(q^{p^{j}}-1)/r)}\right)\\ &=\gcd(r,n/p)=1.\end{split}$ (6.11)
Now, applying Lemma 2.3 with (6.8), (6.10) and (6.11) and using the definition of the Möbius function, we obtain
$N_{1}$ $=0,$
$N_{p}$ $=\frac{\gcd(n,(q^{p}-1)/r)-0}{p}=\frac{p-0}{p}=1,$
$N_{p^{b}}$ $=\Big(\sum_{j=0}^{b}\mu(p^{b-j})\gcd(n,(q^{p^{j}}-1)/r)\Big)/p^{b}$
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}$
$=\frac{n-p}{p^{b}}.$
Moreover, noticing $pN_{p}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have
$\{l_{i}:\,i\in Z_{n,r}\}=\{p,p^{b}\}.$
Then, the desired result follows directly from Theorems 3.1 and $3.$ ∎
We remark that in Theorems 6.1, 6.2 and 6.3, if furthermore $r\leq 2$, then by Theorem 3.3, we can get a better lower bound on minimum distance for the second family of constacyclic codes.
In Table LABEL:tab:pb+1c, we present some codes constructed in Section 6 having good parameters.
Table 6.1: Codes in Section 6
$q$ $p$ $b$ $r$ Lower Bose Parameter Optimality
bound distance
22 22 33 11 22 33 $17,8,6]^{*}$ optimal
22 22 33 11 22 33 $[17,9,5]$ optimal
22 33 22 11 88 $12$ $73,36,14]^{*}$ $d_{\text{best}}=16$
22 33 22 11 88 $11$ $[73,37,13]$ $d_{\text{best}}=14$
33 22 22 11 22 33 $10,5,4]^{*}$ optimal
33 22 22 22 22 66 $10,4,6]^{*}$ optimal
33 22 22 22 11 33 $[10,6,4]$ optimal
33 22 33 22 88 99 $82,40,18]^{*}$ $d_{\text{best}}=21$
44 22 22 11 33 44 $17,8,6]^{*}$ $d_{\text{best}}=8$
44 22 22 33 33 77 $17,8,8]^{*}$ best-known
44 22 22 33 22 77 $[17,9,7]$ best-known
55 22 22 22 44 55 $26,13,8]^{*}$ $d_{\text{best}}=10$
55 22 22 44 44 99 $26,12,10]^{*}$ $d_{\text{best}}=11$
55 22 22 44 33 44 $[26,14,8]$ $d_{\text{best}}=9$
77 22 22 11 77 $10$ $50,25,18]^{*}$ best-known
77 22 22 22 77 99 $[50,26,14]$ $d_{\text{best}}=15$
88 22 22 11 $10$ $22$ $65,32,22]^{*}$ best-known
7 Constacyclic codes with $n=\Phi_{p_{1}p_{2}}(q)$
Recall that $q$ is a prime power, and $r\mid q-1$. Besides, we assume that $p_{1},p_{2}$ are two distinct primes with $p_{1}<p_{2}$. In the following, we construct some infinite families of constacyclic codes with
$n=\Phi_{p_{1}p_{2}}(q)=\frac{(q^{p_{1}p_{2}}-1)(q-1)}{(q^{p_{1}}-1)(q^{p_{2}}-1)}=\frac{q^{p_{1}p_{2}}-1}{\mathrm{lcm}(q^{p_{1}}-1,q^{p_{2}}-1)}.$
Moreover, we only consider the case when the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$, has exactly two elements, and then we use Theorem 3.1 to construct constacyclic codes having good lower bounds for their minimum distances.
Theorem 7.1.
Let $p_{1},p_{2}$ be two distinct primes with $p_{1}<p_{2}$, and $n=\frac{q^{p_{1}p_{2}}-1}{\mathrm{lcm}(q^{p_{1}}-1,q^{p_{2}}-1)}$. Assume $p_{2}\nmid(q^{p_{1}}-1)/r$. Then, an infinite family of constacyclic $n,n-1-p_{1}p_{2}\lceil(n-1)/(2p_{1}p_{2})\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2p_{1}p_{2})\rceil}{q-1}\right\rceil+\Lambda_{p_{1}p_{2}}]$ codes is constructed, where $\Lambda_{p_{1}p_{2}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p_{1}p_{2}(q-1)\geq qr$, an infinite family of constacyclic $n,n-p_{1}p_{2}\lfloor(n-1)/(2p_{1}p_{2})\rfloor,\geq\lfloor\frac{q(n-1)}{2p_{1}p_{2}(q-1)}\rfloor]$ codes is constructed.
Proof.
First, we want to determine $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. Since $r\mid q-1$, we write
$nr=\frac{q^{p_{1}p_{2}}-1}{\mathrm{lcm}(q^{p_{1}}-1,q^{p_{2}}-1)/r},$ (7.1)
where the denomenator is a positive integer. Combining (7.1) with Lemma 2.1, we have
$l_{i}\mid p_{1}p_{2}\quad\textrm{for}\quad i\in Z_{n,r}.$
By Lemma 2.16, we have
$\gcd(n,q-1)=1.$ (7.2)
This, together with $r\mid q-1$, implies that for any integer $j\geq 1$, we have
$\gcd\left(r,\frac{nr}{\gcd(nr,q^{j}-1)}\right)=\gcd\left(r,\frac{n}{\gcd(n,(q^{j}-1)/r)}\right)=1.$ (7.3)
Notice that
$\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1}$ $=q^{p_{1}(p_{2}-1)}+q^{p_{1}(p_{2}-2)}+\cdots+q^{p_{1}}+1$
$\equiv 1+1+\cdots+1+1\equiv p_{2}\pmod{(q^{p_{1}}-1)/r}$.
Then, we get
$\gcd\left(\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1},\frac{q^{p_{1}}-1}{r}\right)=\gcd\left(p_{2},\frac{q^{p_{1}}-1}{r}\right).$
So, noticing $\frac{q^{p_{1}p_{2}}-1}{q^{p_{1}}-1}=n\cdot\frac{q^{p_{2}}-1}{q-1}$ and $\gcd(p_{2},q^{p_{2}}-1)=1$, we obtain
$\gcd\left(n,\frac{q^{p_{1}}-1}{r}\right)=\gcd\left(p_{2},\frac{q^{p_{1}}-1}{r}\right).$ (7.4)
Combining this with the assumption $p_{2}\nmid(q^{p_{1}}-1)/r$, we have
$\gcd\left(n,\frac{q^{p_{1}}-1}{r}\right)=1.$ (7.5)
Now, applying Lemma 2.3 with (7.2), (7.3) and (7.5), we obtain
$N_{1}$ $=\gcd(n,(q-1)/r)=1,$
$N_{p_{1}}$ $=\left(\gcd\left(n,\frac{q^{p_{1}}-1}{r}\right)-N_{1}\right)/p_{1}=(1-1)/p_{1}=0.$
Next, we want to compute $N_{p_{2}}$. Similarly as (7.4), we have
$\gcd\left(n,\frac{q^{p_{2}}-1}{r}\right)=\gcd\left(p_{1},\frac{q^{p_{2}}-1}{r}\right).$
Combining this with $N_{1}=1$ and the assumption $p_{1}<p_{2}$, we get
$N_{p_{2}}=\left(\gcd\left(n,\frac{q^{p_{2}}-1}{r}\right)-N_{1}\right)/p_{2}\leq(p_{1}-1)/p_{2}<1.$
Thus, we have $N_{p_{2}}=0$.
In addition, we directly have
$N_{p_{1}p_{2}}=\frac{n-N_{1}-p_{1}N_{p_{1}}-p_{2}N_{p_{2}}}{p_{1}p_{2}}=\frac{n-1}{p_{1}p_{2}}.$
Hence, we have
$\{l_{i}:\,i\in Z_{n,r}\}=\{1,p_{1}p_{2}\}.$
Then, the desired results follow directly from Theorems 3.1 and $3.$ ∎
We remark that the condition $p_{2}\nmid(q^{p_{1}}-1)/r$ in Theorem 7.1 holds when $p_{1}\nmid p_{2}-1$. Indeed, if $p_{1}\nmid p_{2}-1$, then $q^{p_{1}}\not\equiv 1\pmod{p_{2}}$, and so $p_{2}\nmid q^{p_{1}}-1$.
We also remark that from (7.4) and the computation of $N_{p_{1}}$, one can see that $N_{p_{1}}=0$ if and only if $p_{2}\nmid(q^{p_{1}}-1)/r$. This means that the set $\{l_{i}:\,i\in Z_{n,r}\}$ has exactly two elements if and only if $p_{2}\nmid(q^{p_{1}}-1)/r$.
In addition, in Theorem 5.1, if furthermore $r\leq 2$, then by Theorem 3.3, we know that for the second family of constacyclic codes, the lower bound $\lfloor\frac{q(n-1)}{2p_{1}p_{2}(q-1)}\rfloor$ can be replaced by
$\left\lceil\frac{q\lfloor(n-1)/(2p_{1}p_{2})\rfloor}{q-1}\right\rceil+\Lambda_{p_{1}p_{2}}^{\prime},$
where $\Lambda_{p_{1}p_{2}}^{\prime}$ has been defined in Theorem 3.3.
In Table LABEL:tab:thmp1p2f, we present some codes constructed in Theorem 7.1 having good parameters.
Table 7.1: Codes in Theorem 7.1
$q$ $p_{1}$ $p_{2}$ $r$ Lower Bose Parameter Optimality
bound distance
22 22 77 11 44 77 $43,14,14]^{*}$ optimal
22 22 77 11 33 33 $[43,29,6]$ optimal
33 22 55 11 55 66 $61,30,15]^{*}$ $d_{\text{best}}=18$
33 22 55 11 44 55 $[61,31,14]$ $d_{\text{best}}=16$
44 22 33 33 22 33 $13,6,6]^{*}$ optimal
44 22 33 33 11 33 $[13,7,5]$ optimal
77 22 33 11 55 66 $43,18,16]^{*}$ $d_{\text{best}}=18$
77 22 33 22 55 88 $43,18,17]^{*}$ $d_{\text{best}}=18$
77 22 33 22 44 55 $[43,25,10]$ $d_{\text{best}}=12$
77 22 33 33 55 88 $43,18,17]^{*}$ $d_{\text{best}}=18$
77 22 33 33 44 55 $[43,25,11]$ $d_{\text{best}}=12$
77 22 33 66 55 77 $43,18,15]^{*}$ $d_{\text{best}}=18$
77 22 33 66 44 66 $[43,25,11]$ $d_{\text{best}}=12$